# Towards Efficient Medical Reasoning with Minimal Fine-Tuning Data

### v2

@@ Abstract
Supervised Fine-Tuning (SFT) plays a pivotal role in adapting Large Language Models (LLMs) to specialized domains such as medical reasoning. However, existing SFT practices often rely on unfiltered datasets that contain redundant and low-quality samples, leading to substantial computational costs and suboptimal performance. Although existing methods attempt to alleviate this problem by selecting data based on sample difficulty, defined by knowledge and reasoning complexity, they overlook each sample’s optimization utility reflected in its gradient. Interestingly, we find that gradient-based influence alone favors easy-to-optimize samples that cause large parameter shifts but lack deep reasoning chains, while difficulty alone selects noisy or overly complex cases that fail to guide stable optimization. Based on this observation, we propose a data selection strategy, Difficulty-Influence Quadrant (DIQ), which prioritizes samples in the “high-difficulty–high-influence” quadrant to balance complex clinical reasoning with substantial gradient influence, enabling efficient medical reasoning with minimal fine-tuning data. Furthermore, Human and LLM-as-a-judge evaluations show that DIQ-selected subsets demonstrate higher data quality and generate clinical reasoning that is more aligned with expert practices in differential diagnosis, safety check, and evidence citation, as DIQ emphasizes samples that foster expert-like reasoning patterns. Extensive experiments on medical reasoning benchmarks demonstrate that DIQ enables models fine-tuned on only 1% of selected data to match full-dataset performance, while using 10% consistently outperforms baseline methods, highlighting the superiority of principled data selection over brute-force scaling.
@@ 1 Introduction
Large Language Models (LLMs) [7, 24] have achieved notable success in reasoning-intensive tasks such as mathematics and programming [7, 21]. Inspired by this progress, the medical community is exploring LLM for high-stakes scenarios such as clinical diagnosis and treatment planning, and the prospect of LLM that emulates clinician cognitive processes is transforming the landscape of healthcare services [13, 23, 29]. However, current models still face significant challenges in synthesizing incomplete and ambiguous clinical information into reliable decisions [12, 22].
@@
Figure 1: (a) The FineMed dataset distributed by difficulty and influence scores, with data points colored by quadrant. (b) For each quadrant, the bar shows the intrinsic reasoning quality of the data and the resulting downstream performance of a Qwen3-8B model fine-tuned on that subset.
@@
Supervised Fine-Tuning (SFT) is a prevailing practice for adapting LLMs to medical domains. Following the assumption that larger datasets improve performance, previous works explored scaling up medical datasets with tens of thousands of Chain-of-Thought (CoT) examples [34, 20]. For instance, ReasonMed employs multi-agent verification to curate 370k instances [20]. However, reliance on unfiltered data with redundant and low-value samples incurs substantial computational costs and inefficient training. Early data selection efforts have relied on coarse, difficulty-based heuristics, such as removing easy samples or retaining only those unsolved after multiple trials [2, 28]. While more recent studies refine this by decoupling difficulty into Knowledge and Reasoning complexity [22, 9], these approaches are still limited. Critically, they remain one-dimensional, evaluating samples in isolation while overlooking their optimization utility as reflected by the training gradient. Consequently, the precise impact of difficult samples in training, encompassing both their benefits and drawbacks, is not yet fully understood.
@@
To investigate the interplay between sample difficulty and influence, we conducted a pilot experiment on the FineMed dataset [33]. We partitioned the data into four quadrants based on these two metrics (Fig. 1 (a)) and then evaluated each quadrant for model performance (via fine-tuning on them) and reasoning quality, as assessed by Gemini-2.5-Pro, as detailed in Fig. 1 (b) and App. A. The results reveal a critical tension. Models fine-tuned on high-influence, low-difficulty $\mathcal{Q}_{2}$ data consistently outperformed those trained on low-influence, high-difficulty data $\mathcal{Q}_{3}$, even though the samples in $\mathcal{Q}_{2}$ are of lower reasoning quality. We found this occurs because $\mathcal{Q}_{2}$ samples, while easy to optimize, possess shallower reasoning chains. Conversely, $\mathcal{Q}_{3}$ samples are reasoning-intensive but provide weak gradient signals, leading to unstable training and lower downstream performance. These findings expose a fundamental flaw in one-dimensional selection: prioritizing influence alone favors simplistic samples, while prioritizing difficulty alone selects for noisy, hard-to-learn cases.
@@
In this paper, we propose a simple yet effective data selection strategy, Difficulty-Influence Quadrant (DIQ), which prioritizes samples in the high-difficulty–high-influence quadrant to balance complex clinical reasoning with substantial influence score, to enable efficient medical reasoning with minimal data. Specifically, the Difficulty score for each medical sample is obtained from a classifier that is constructed by fine-tuning BiomedBERT [6] on medical questions collected from multiple medical QA datasets and assesses the complexity of knowledge and reasoning on a 5-point Likert scale. Meanwhile, the Influence score quantifies the expected impact of each instance on model improvement, efficiently calculated by aggregating first-order gradient dot products between training and validation samples across tasks and epochs. This lightweight approach enables scalable influence approximation while avoiding the high computational cost of traditional methods. Based on these scores, our selection strategy proceeds in the order of $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, $\mathcal{Q}_{3}$ and $\mathcal{Q}_{4}$ as defined in Fig. 1 (a), to ensure that the selected subset maximally balances the complexity of reasoning and the utility of training.
@@
To assess the effectiveness of DIQ, we perform human and LLM-as-a-judge evaluations on DIQ-selected subsets, showing that they exhibit higher data quality and generate clinical reasoning closely aligned with expert practices in differential diagnosis, safety check, and evidence citation. These results demonstrate that DIQ selects samples that foster expert-like reasoning patterns. Extensive experiments on medical reasoning benchmarks demonstrate that DIQ enables models fine-tuned on only 1% of selected data to match full-dataset performance, while 10% consistently outperforms baseline methods, highlighting the superiority of principled data selection over brute-force scaling. Our contributions can be summarized as follows:
@@
  • We propose DIQ, a principled data selection framework for medical SFT that jointly measures sample difficulty and optimization influence, enabling more efficient medical reasoning with minimal fine-tuning data.
@@ 2 Related Work
The construction of high-quality medical reasoning datasets has emerged as a central focus, with methodologies divided into two distinct paradigms. The predominant approach involves creating large-scale datasets through synthetic data generation using powerful foundation models like GPT-4o within multi-agent verification pipelines (e.g., UltraMedical [34], ReasonMed [20]), or by leveraging medical knowledge graphs and multi-stage training procedures to embed clinical validity and complex reasoning pathways (e.g., MedReason [28], FineMedLM-o1 [33], HuatuoGPT-o1 [2]). Despite driving recent progress, large synthetic datasets remain costly and labor-intensive to curate consistently. In contrast, a compelling line of research [9, 35, 32] posits that for base models with extensive domain knowledge, complex reasoning capabilities can be effectively elicited with remarkably few high-quality examples. Inspired by this less is more premise, our work investigates data-efficient fine-tuning for medical reasoning.
@@ 3.1 Overview
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\sum_{z\in\mathcal{D}}\ell(z;\boldsymbol{\theta}),$ (1)
where $\ell$ denotes the token-level cross-entropy loss. Our goal is to select a subset $\mathcal{S}\subset\mathcal{D}$ with the pre-defined retention ratio $r\in(0,1)$ such that SFT on $\mathcal{S}$ maximizes downstream performance $\mathcal{M}(\boldsymbol{\theta}^{\mathcal{S}})$.
@@ 3.3 Influence via Gradient Dot-Product
Intuitively, $\mathrm{Dot}(z)$ measures the alignment between the training gradient of $z$ and the average validation gradient. A larger positive value implies a stronger expected decrease of the average validation loss after updating on $z$; a negative value suggests potential harm.
@@
$\Delta\bar{\ell}_{\text{val}}$ $=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{z^{\prime}}\big(\ell(z^{\prime};\boldsymbol{\theta}^{+})-\ell(z^{\prime};\boldsymbol{\theta})\big)$ (6)
@@
In practice, we first compute $\bar{g}_{\text{val}}$ in Eq. (3) with a single backward pass per validation sample. Then, for each training sample $z$, we compute one backward pass to obtain $g(z)$ and take the inner product in Eq. (4). This yields an $O(|\mathcal{D}_{\text{val}}|+|\mathcal{D}|)$ backward complexity and requires no Hessian or Hessian-vector products. Moreover, we apply a Johnson–Lindenstrauss guaranteed random projection to restrict gradient to a lower-dimension subspace (the default dimension is set as 4096), which preserves ranking well and reduces computational cost enormously.
@@ 3.4 Quadrant-Based Data Selection
Given the two score dimensions $\big(D(z),\,\mathrm{Dot}(z)\big)$, we partition $\mathcal{D}$ into four quadrants using data-adaptive thresholds. Let $\tau_{d}$ denote the difficulty threshold (a percentile of $\{D(z)\}$), and let $m_{\text{dot}}$ denote the median of $\{\mathrm{Dot}(z)\}$, the four quadrants can be defined as
@@ 4.2 Main Results
To isolate the effect of DIQ when complex intermediate reasoning is absent, we construct MedReason-QA by removing CoT traces from MedReason [28] while keeping the original splits and answer keys. As shown in Fig. 3, DIQ consistently outperforms random selection at both 1% and 10% data budgets for Llama3.1-8B-Instruct and Qwen3-8B, with averages computed over five runs. This indicates that DIQ does not rely on explicit reasoning traces: even under QA-only supervision, where labels are short and signal-to-noise is lower, DIQ still identifies high–information density samples and preserves coverage of rare and clinically important cases. Practically, when CoT annotations are unavailable or costly, DIQ serves as a drop-in data selector for medical QA datasets, delivering robust gains under tight budgets. Full results are reported in App. G.2.
@@
Data DDx Safety Check Evidence Citation
@@
    All three evaluation metrics are rated on a 5-point scale (1-5).
Table 2: Clinical value comparison: DIQ-1% Data vs. remainder, and DIQ-1% model vs. full-dataset model.
@@ 5.3 Effect of Validation Set Size
DIQ relies on a held-out validation set to estimate example influence and calibrate difficulty, so its size trades off estimator stability against compute. As shown in Table 3 (Llama3.1-8B-Instruct on Huatuo at a 1% keeping ratio), enlarging the validation split from 90 to 450 yields steady overall gains (AvgA: 42.04 $\rightarrow$ 43.89, +1.85). Improvements are not linear: most of the benefit is realized by 360 examples (AvgA=43.80, +1.76 over 90), after which returns are marginal (+0.09 at 450). By dimension, AvgS grows nearly monotonically (56.03 $\rightarrow$ 57.76), whereas AvgC is mildly non-monotonic—peaking at 360 (36.99) and remaining numerically comparable at 450 (36.95, $\Delta$=0.04). These trends suggest that moderate validation sizes already stabilize DIQ’s influence rankings while avoiding the extra compute of very large splits; in practice, 360–450 examples provide a strong quality–cost trade-off under this budget.
@@ 6 Conclusion
In this paper, we introduce DIQ, a data selection framework that identifies compact yet high‑value training subsets for medical reasoning by jointly considering sample difficulty and model‑dependent influence via a simple, Hessian‑free gradient inner product (Dot). DIQ maps each example into the difficulty–influence plane and prioritizes the high‑difficulty/high‑influence region, yielding a principled, model‑aware curriculum. Across nine downstream tasks, fine‑tuning on only 1–10% of DIQ‑selected data matches or surpasses training on the full dataset, consistently preserving or improving accuracy while substantially reducing the training data footprint and computational cost.
@@ C.2 Difficulty Classifier Training
We evaluated three lightweight BERT-style models for predicting medical difficulty: BiomedBERT, ClinicalBERT [25], and ModernBERT [27]. As shown in Table 6, BiomedBERT consistently outperformed the other models and was therefore selected as our difficulty classifier.

### v3

@@ Abstract
Supervised Fine-Tuning (SFT) of the language backbone plays a pivotal role in adapting Vision-Language Models (VLMs) to specialized domains such as medical reasoning. However, existing SFT practices often rely on unfiltered textual datasets that contain redundant and low-quality samples, leading to substantial computational costs and suboptimal performance in complex clinical scenarios. Although existing methods attempt to alleviate this problem by selecting data based on sample difficulty, defined by knowledge and reasoning complexity, they overlook each sample’s optimization utility reflected in its gradient. Interestingly, we find that gradient-based influence alone favors easy-to-optimize samples that cause large parameter shifts but lack deep reasoning chains, while difficulty alone selects noisy or overly complex textual cases that fail to guide stable optimization. Based on this observation, we propose a data selection strategy, Difficulty-Influence Quadrant (DIQ), which prioritizes samples in the ”high-difficulty-high-influence” quadrant to balance complex clinical reasoning with substantial gradient influence. This enables efficient medical reasoning for VLMs with minimal fine-tuning data. Furthermore, Human and LLM-as-a-judge evaluations show that DIQ-selected subsets demonstrate higher data quality and generate clinical reasoning that is more aligned with expert practices in differential diagnosis, safety check, and evidence citation, as DIQ emphasizes samples that foster expert-like reasoning patterns. Extensive experiments on medical reasoning benchmarks demonstrate that DIQ enables VLM backbones fine-tuned on only 1% of selected data to match full-dataset performance, while using 10% consistently outperforms baseline methods, highlighting the superiority of principled data selection over brute-force scaling. The code is available at https://github.com/mihara-bot/DIQ.
@@ 1 Introduction
Large Language Models (LLMs) [7, 26], which serve as the core reasoning engines for modern Vision-Language Models (VLMs), have achieved notable success in reasoning-intensive tasks such as mathematics and programming [7, 23]. Inspired by this progress, the medical community is exploring these foundation models for high-stakes scenarios such as clinical diagnosis and treatment planning, and the prospect of language backbones that emulates clinician cognitive processes is transforming the landscape of healthcare services, especially in multimodal applications [15, 25, 30]. However, current models still face significant challenges in synthesizing incomplete and ambiguous clinical information (often derived from complex visual findings) into reliable decisions [14, 24].
@@
Figure 1: (a) The FineMed dataset distributed by difficulty and dot influence scores, with data points colored by quadrant. (b) For each quadrant, the bar shows the intrinsic reasoning quality of the data and the resulting downstream performance of a Qwen3-8B model fine-tuned on the corresponding data subset.
@@
Supervised Fine-Tuning (SFT) is a prevailing practice for adapting LLM backbonesto medical domains. Following the assumption that larger datasets improve performance, previous works explored scaling up medical datasets with tens of thousands of Chain-of-Thought (CoT) examples [35, 22]. For instance, ReasonMed employs multi-agent verification to curate 370k instances [22]. However, reliance on unfiltered data with redundant and low-value samples incurs substantial computational costs and inefficient training. Early data selection efforts have relied on coarse, difficulty-based heuristics, such as removing easy samples or retaining only those unsolved after multiple trials [2, 29]. While more recent studies refine this by decoupling difficulty into Knowledge and Reasoning complexity [24, 10], these approaches are still limited. Critically, they remain one-dimensional, evaluating samples in isolation while overlooking their optimization utility [12, 19] as reflected by the training gradient. Consequently, the precise impact of difficult samples in training, encompassing both their benefits and drawbacks, is not yet fully understood.
@@
To investigate the interplay between sample difficulty and influence, we conducted a pilot experiment on the FineMed dataset [34]. We partitioned the data into four quadrants based on these two metrics (Fig. 1 (a)) and then evaluated each quadrant for model performance (via fine-tuning on them) and reasoning quality, as assessed by Gemini-2.5-Pro [3], as detailed in Fig. 1 (b) and App. A. The results reveal a critical tension. Models fine-tuned on high-influence, low-difficulty $\mathcal{Q}_{2}$ data consistently outperformed those trained on low-influence, high-difficulty data $\mathcal{Q}_{3}$, even though the samples in $\mathcal{Q}_{2}$ are of lower reasoning quality. We found this occurs because $\mathcal{Q}_{2}$ samples, while easy to optimize, possess shallower reasoning chains. Conversely, $\mathcal{Q}_{3}$ samples are reasoning-intensive but provide weak gradient signals, leading to unstable training and lower downstream performance. These findings expose an unavoidable flaw in one-dimensional selection: prioritizing influence alone favors simplistic samples, while prioritizing difficulty alone selects for noisy, hard-to-learn cases.
@@
In this paper, we propose a simple yet effective data selection strategy for textual medical reasoning, Difficulty-Influence Quadrant (DIQ), which prioritizes samples in the high-difficulty-high-influence quadrant to balance complex clinical reasoning with substantial influence score, to enable efficient medical reasoning with minimal data. Specifically, the Difficulty for each medical sample is obtained from a classifier constructed by fine-tuning BiomedBERT [6] on medical questions collected from multiple medical QA datasets and assesses the complexity of knowledge and reasoning on a 5-point Likert scale. Meanwhile, the Influence quantifies the expected impact of each instance on model improvement, efficiently calculated by aggregating first-order gradient dot products between training and validation samples across tasks and epochs. This lightweight approach enables scalable influence approximation while avoiding the high computational cost of traditional methods. Based on these two scores, our selection strategy proceeds in the order of $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, $\mathcal{Q}_{3}$ and $\mathcal{Q}_{4}$ as defined in Fig. 1 (a), to ensure that the selected subset maximally balances the complexity of reasoning and the utility of training.
@@
To assess the effectiveness of DIQ, we perform human and LLM-as-a-judge evaluations on DIQ-selected subsets, showing that they exhibit higher data quality and generate clinical reasoning closely aligned with expert practices in differential diagnosis, safety check, and evidence citation (Tab. 2). These results demonstrate that DIQ selects samples that foster expert-like reasoning patterns. Extensive experiments on medical reasoning benchmarks demonstrate that DIQ enables models fine-tuned on only 1% of selected data to match full-dataset performance, while 10% consistently outperforms baseline methods, highlighting the superiority of principled data selection over brute-force scaling. Our contributions can be summarized as follows:
@@
  • We propose DIQ, a principled data selection framework for medical SFT that jointly measures sample difficulty and optimization influence, enabling more efficient medical reasoning with fewer fine-tuning data, thereby building robust language foundations for medical VLMs.
@@ 2 Related Work
The construction of high-quality medical reasoning datasets has emerged as a central focus, with methodologies divided into two distinct paradigms. The predominant approach involves creating large-scale datasets through synthetic data generation using powerful foundation models like GPT-4o within multi-agent verification pipelines (e.g., UltraMedical [35], ReasonMed [22]), or by leveraging medical knowledge graphs and multi-stage training procedures to embed clinical validity and complex reasoning pathways (e.g., MedReason [29], FineMedLM-o1 [34], HuatuoGPT-o1 [2]). Despite driving recent progress, large synthetic datasets remain costly and labor-intensive to curate consistently. In contrast, a compelling line of research [10, 36, 33] posits that for base models with extensive domain knowledge, complex reasoning capabilities can be effectively elicited with remarkably few high-quality examples. Inspired by this less is more premise, our work investigates data-efficient fine-tuning for medical reasoning. We demonstrate that strategically selecting a small subset from existing medical SFT datasets can achieve performance comparable or even superior to full-dataset fine-tuning.
@@ 3.1 Overview
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathbb{E}_{(q,a)\sim\mathcal{D}}[\ell(q,a;\boldsymbol{\theta})],$ (1)
where $\ell$ denotes the token-level cross-entropy loss. Our goal is to select a subset $\mathcal{S}\subset\mathcal{D}$ with the pre-defined retention ratio $r\in(0,1)$ such that SFT on $\mathcal{S}$ maximizes downstream performance of the model.
@@ 3.3 Dot-Product Influence
Intuitively, $\mathrm{Dot}(z)$ measures the alignment between the training gradient of $z$ and the average validation gradient. A larger positive value implies a stronger expected decrease of the average validation loss after updating on $z$ while a negative value suggests potential harm.
@@
$\Delta\bar{\ell}_{\text{val}}$ $=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{z^{\prime}\in\mathcal{D}_{\text{val}}}\big(\ell(z^{\prime};\boldsymbol{\theta}^{+})-\ell(z^{\prime};\boldsymbol{\theta})\big)$ (6)
@@
In practice, we first compute $\bar{g}_{\text{val}}$ in Eq. (3) with a single backward pass per validation sample. Then, for each training sample $z$, we compute one backward pass to obtain $g(z)$ and take the inner product in Eq. (4). This yields an $O(|\mathcal{D}_{\text{val}}|+|\mathcal{D}|)$ backward complexity and requires no Hessian or Hessian-vector products. Moreover, we apply a Johnson-Lindenstrauss guaranteed random projection to restrict gradient to a lower-dimension subspace (the default dimension is set as 4096), which preserves ranking well and reduces computational cost enormously.
@@ 3.4 Quadrant-Based Data Selection
Having computed the two score dimensions $D(z),\,\mathrm{Dot}(z)$, we partition $\mathcal{D}$ into four quadrants using data-adaptive thresholds. Let $\tau_{d}$ denote the difficulty threshold (a percentile of $\{D(z)\}$), and let $m_{\text{dot}}$ denote the median of $\{\mathrm{Dot}(z)\}$, the four quadrants can be defined as
@@ 4.2 Main Results
To isolate the effect of DIQ when complex intermediate reasoning is absent, we construct MedReason-QA by removing CoT traces from MedReason [29] while keeping the original splits and answer keys. As shown in Fig. 3, DIQ consistently outperforms random selection at both 1% and 10% data budgets for Llama3.1-8B-Instruct and Qwen3-8B, with averages computed over five runs. This indicates that DIQ does not rely on explicit reasoning traces: even under QA-only supervision, where labels are short and signal-to-noise is lower, DIQ still identifies high-information density samples and preserves coverage of rare and clinically important cases. Practically, when CoT annotations are unavailable or costly, DIQ serves as a drop-in data selector for medical QA datasets, delivering robust gains under tight budgets. Full results are reported in App. G.2.
@@
Data DDx SC EC
@@
Table 2: Clinical value comparison: DIQ-1% Data vs. remainder, and DIQ-1% model vs. full-dataset model. All three evaluation metrics are rated on a 5-point scale. Abbreviations: DDx = Differential Diagnosis, SC = Safety Check, EC = Evidence Citation.
@@ 5.3 Effect of Validation Set Size
DIQ relies on a held-out validation set to estimate example influence and calibrate difficulty, so its size trades off estimator stability against compute. As shown in Table 3 (Llama3.1-8B-Instruct on Huatuo at a 1% keeping ratio), enlarging the validation split from 90 to 450 yields steady overall gains (AvgA: 42.04 $\rightarrow$ 43.89, +1.85). Improvements are not linear: most of the benefit is realized by 360 examples (AvgA=43.80, +1.76 over 90), after which returns are marginal (+0.09 at 450). By dimension, AvgS grows nearly monotonically (56.03 $\rightarrow$ 57.76), whereas AvgC is mildly non-monotonic—peaking at 360 (36.99) and remaining numerically comparable at 450 (36.95, $\Delta$=0.04). These trends suggest that moderate validation sizes already stabilize DIQ’s influence rankings while avoiding the extra compute of very large splits; in practice, 360–450 examples provide a strong quality-cost trade-off under this budget.
@@ 6 Conclusion
In this paper, we introduce DIQ, a data selection framework that identifies compact yet high-value training subsets for medical reasoning by jointly considering sample difficulty and model-dependent influence via a simple gradient inner product (Dot). DIQ maps each example into the difficulty-influence plane and prioritizes the high-difficulty/influence region, yielding a principled, model-aware curriculum. Across nine downstream tasks, fine-tuning on only 1–10% of DIQ-selected data matches or surpasses training on the full dataset, consistently preserving or improving accuracy while substantially reducing the training data footprint and computational cost.
@@
Acknowledgement
This work is supported by the National Key Research and Development Program of China under Grant 2024YFC3308500.
@@ C.2 Difficulty Classifier Training
We evaluated three lightweight BERT-style models for predicting medical difficulty: BiomedBERT [6], ClinicalBERT [9], and ModernBERT [28]. As shown in Table 6, BiomedBERT consistently outperformed the other models and was therefore selected as our difficulty classifier.